# Parameter-Efficient Fine-Tuning in Spectral Domain for Point Cloud Learning

### v1

@@ 6.5 Compared with PEFT methods from other domains
The provided Tab. VIII presents a comparative analysis of various PEFT methods originating from natural language processing (NLP) and 2D vision. The results clearly show that although PEFT methods in other areas can be adapted for 3D tasks, their performance falls short compared to methods specifically designed for 3D vision. For instance, the best-performing PEFT methods from NLP and 2D vision achieve overall accuracies of 83.93% and 83.66%, respectively, leaving a significant performance gap compared to fully fine-tuning, which achieves 85.18%. This indicates the need for 3D-specific adaptations or entirely new methods tailored to 3D data. The existing point cloud PEFT methods, IDPT [88] and DAPT [99], show better performance with accuracies of 84.94% and 85.08%, respectively. However, their trainable parameters are noticeably larger than NLP and 2D counterparts and can not generalized well in different point cloud datasets or pre-trained models (see Sec. 6.1.1). In contrast, the proposed PointGST outperforms all other methods with an accuracy of 85.29%, surpassing both IDPT and DAPT, with only 0.6M trainable parameters.
TABLE VIII: Comparisons of Parameter-Efficient Fine-Tuning (PEFT) methods from NLP and 2D Vision on the hardest variant of ScanObjectNN [68]. Overall accuracy (%) without voting is reported. Params. represents the trainable parameters.
@@ 6.6 Visualization Results
6.7 Ablation Study
Following previous methods [99, 88], we conduct ablation studies on the challenging ScanObjectNN PB_T50_RS dataset. Unless otherwise specified, we choose the Point-MAE [47] and PointGPT-L [4] as the baseline, and all experiments are performed without the voting strategy.
6.7.1 The effect of each component
We systematically evaluate the effect of each component within our PointGST by analyzing four key aspects: global spectral, local spectral, residual connections, and spectral inversion. The results are detailed in Tab. IX.
Taking the Point-MAE baseline as an example, without any of our components, the linear probing setting achieves only 75.99% overall accuracy (OA), significantly lower than the fully fine-tuning counterpart by 9.19%. When incorporating either global or local spectral components individually, we observe a marked improvement of 6.00% in OA for both, while introducing only 0.35M additional trainable parameters. Further performance gains are seen when integrating residual connections and applying the inverse Graph Fourier Transform (iGFT) to project back into the spatial domain, aligning with the output of the transformer layer. Ultimately, the combination of all components leads to a significant increase in accuracy to 85.29% OA. This outcome underscores the complementary nature of global and local spectral representations, which can effectively de-correlate the inner confusion of point tokens and provide task-related intrinsic information to guide the fine-tuning phase.
To further demonstrate the effectiveness of fine-tuning in the spectral domain, we apply different tokens in the proposed Point Cloud Spectral Adapter (Fig. 2(b)), including the default point tokens $\boldsymbol{T}_{s}$, initializing a learnable matrix to act on $\boldsymbol{T}_{s}$, and our method of transferring the $\boldsymbol{T}_{s}$ (serving as graph signal) to the spectral tokens. As shown in Fig. 7, the proposed paradigm achieves better performance compared with the other two settings, demonstrating the importance of fine-tuning in the spectral domain.
TABLE IX: The effect of each component of our PointGST. The trainable parameters (Params.) and the overall accuracy (%) on the hardest variant of ScanObjectNN [68] are reported. The results before and after the slash “/” indicate the results when using Point-MAE and PointGPT-L as the baselines, respectively.
@@ 6.7.1 The effect of each component
      ✓       -       -       -       0.62/2.43       81.99/93.88
      ✓       -       ✓       -       0.62/2.43       83.03/94.27
      ✓       -       -       ✓       0.62/2.43       84.07/94.48
      ✓       -       ✓       ✓       0.62/2.43       84.52/94.69
      -       ✓       -       -       0.62/2.43       81.99/94.17
      -       ✓       ✓       -       0.62/2.43       83.38/94.45
      -       ✓       -       ✓       0.62/2.43       83.69/94.55
      -       ✓       ✓       ✓       0.62/2.43       83.83/94.62
      ✓       ✓       ✓       ✓       0.62/2.43       85.29/94.83
6.7.2 Ablation study on the hyper-parameters and designs choices of PCSA
The effect of dimension $r$. The PCSA aims to decrease the trainable parameters during fine-tuning, and the dimension $r$ is chosen to balance efficiency and performance. As shown in Tab. X(a), as the dimension $r$ increases from 12 to 36, the performance of our method improves, with performance increasing from 83.66 to 85.29. However, further increasing $r$ to 96 slightly hurts the performance, with performance dropping to 84.98. Therefore, we set the dimension $r$ to 36 by default, which reports ideal performance and less trainable parameters.
The effect of scale $s$. The scaling factor $s$ can control the impact of spectral tokens on the task-agnostic tokens. As shown in Tab. X(b), we empirically find that when the value is simply set as $1$ without the need for extensive hyper-parameter tuning, the proposed can achieve satisfactory performance.
The effect of different adaption choices. In our PCSA, we utilize a simple shared linear layer to adapt the global and local spectral tokens, shown in Fig. 2(b). Here, we explore other possible learnable choices, including shared Attention, MLP (two layers), DWConv, linear, and individual linear for global and local spectral. Note that for fair comparisons, we make these options meet similar trainable parameters (about 0.6M). The results presented in Tab. X(c) validate that the used shared linear layer is simple yet effective compared with other learnable architectures.
TABLE X: Ablation study on the Point Cloud Spectral Adapter (PCSA), including dimension $r$, scale $s$, and the effect of different adaption choices. The trainable parameters (Params.) and the overall accuracy (%) on the hardest variant of ScanObjectNN [68] are reported. The results before and after the slash “/” indicate the results when using Point-MAE and PointGPT-L as the baselines, respectively.
@@ 6.7.2 Ablation study on the hyper-parameters and designs choices of PCSA
TABLE XI: The effect of inserted position. The trainable parameters (Params.) and the overall accuracy (%) on the hardest variant of ScanObjectNN [68] are reported. The results before and after the slash “/” indicate the results when using Point-MAE and PointGPT-L as the baselines, respectively.
@@ 6.7.3 The effect of the position of PCSA
We then study the effect of the position of the PCSA. For the inserted layers, we consider the shallow layers and deep layers. Tab. XI(a) clearly demonstrates that the performance of PointGST has a positive relative to the added layers. Taking the Point-MAE baseline as an example, injecting the PCSA into half of the layers will cause about a 2.12%/1.53% performance drop. Therefore, in our method, we inject the proposed point spectral adapter into all transformer layers. In addition, we also explore the effect of inserted modules. There are three potential choices: FFN, Attention, and each entire transformer block. As shown in Tab. XI(b), inserting the PCSA into FFN is able to outperform the Attention and entire transformer block by 1.01% and 1.04% (resp. 0.42%, 0.45%) overall accuracy for Point-MAE (resp. PointGPT-L), respectively.
@@ 6.7.4 Analysis on the point cloud graph construction
The effect of different weighting schemes. To calculate the edge weight between two points in a graph, we employ a data-dependent scaling strategy. Here, we study various weighting schemes, as detailed in Tab. XII(a). The most straightforward weighting scheme, Euclidean distance, yields sub-optimal performance with an overall accuracy of 85.15% and 94.73%. As discussed in Sec. 4.2.1, we contend that the closer two points are, the stronger the weight of their relationship should be. Therefore, we propose a data-dependent scaled distance for edge weight, which outperforms Euclidean distance by 0.14% and 0.10% on the Point-MAE and PointGPT-L baselines. Additionally, we evaluate cosine distance and Gaussian similarity [70] for each pair of key points, finding that our proposed data-dependent scaling strategy yields more stable results.
@@
The effect of different sorting methods and group number $k$. When constructing local sub-graphs, we sort and categorize the sampled key points into $k$ groups. This part evaluates the impact of various sorting methods, including random, k-nearest neighbors (KNN), and space-filling curves such as Hilbert order, transposed Hilbert order, Z-order, and transposed Z-order. The results presented in Tab. XII(b) show that random sorting provides the least favorable results, while methods that preserve spatial locality (e.g., KNN and space-filling curves) considerably enhance the performance metrics. Among them, Trans Z-order sorting stands out as the most effective. KNN exhibits inferior performance, potentially because it does not yield stable results and sometimes aggregates identical points into different groups. In contrast, space-filling curves maintain a consistent scanning pattern, making it easier for models to learn from the data. Besides, as shown in Tab. XII(c), we empirically find that the best performance is achieved when $k$ is set as 4.
TABLE XII: Analysis on the point cloud graph.
(a) The effect of different edge weights.
@@
Original point cloud Euclidean Distance 85.15/94.73
@@
(b) The effect of different sorting methods.
@@
      Hilbert-order       84.00/94.45
      Trans Hilbert-order       84.59/94.34
@@
(c) The effect of group number $k$.
@@
      $2$       84.04/94.45
      $4$       85.29/94.83
      $8$       84.04/94.52
      $16$       84.84/94.55
6.7.5 Analysis on Different Spectral Conversion
We finally investigate the effect of different spectral conversions, including the Fast Fourier Transform (FFT), Discrete Cosine Transform (DCT), and our used Graph Fourier Transform (GFT). As shown in Tab. XIII, the results demonstrate the superiority of the GFT we applied. This can be attributed to the fact that the FFT and DCT are designed for regularly sampled or structured data and assume uniform spacing and regularity in the signals, making them unsuitable for processing non-Euclidean point clouds [25]. In contrast, to accommodate the irregular domain of point clouds, we apply the GFT which is well-suited for such data structures and adapts to specific samples.
TABLE XIII: Analysis on the spectral conversion.

### v2

@@ 6.7 Compared with PEFT Methods from Other Domains
Tab. IX presents a comparative analysis of various PEFT methods originating from natural language processing (NLP) and 2D vision. The results clearly show that although PEFT methods in other areas can be adapted for 3D tasks, their performance falls short compared to methods specifically designed for 3D vision. For instance, the best-performing PEFT methods from NLP and 2D vision achieve overall accuracies of 83.93% and 83.66%, respectively, leaving a significant performance gap compared to fully fine-tuning, which achieves 85.18%. This indicates the need for 3D-specific adaptations or entirely new methods tailored to 3D data. The existing point cloud PEFT methods, IDPT [16] and DAPT [17] show better performance with accuracies of 84.94% and 85.08%, respectively. However, their trainable parameters are noticeably larger than NLP and 2D counterparts and cannot generalize well in different point cloud datasets or pre-trained models (see Sec. 6.1.1). In contrast, the proposed PointGST outperforms all other methods with an accuracy of 85.29%, surpassing IDPT, Point-PEFT [18], and DAPT, with only 0.6M trainable parameters.
TABLE IX: Comparisons of Parameter-Efficient Fine-Tuning (PEFT) methods from NLP and 2D vision on the hardest variant of ScanObjectNN [75]. Overall accuracy (%) without voting is reported. Params. is the trainable parameters.
@@
Figure 7: The comparison of used training GPU memory on NVIDIA 4090.
TABLE X: The comparison of our PointGST and baseline methods in terms of storage usage when fine-tuning various downstream point cloud datasets or tasks.
Point-MAE [8] 22.1M 22.1M 22.1M 22.1M 22.1M 27.1M 27.0M 164.6M
PointGST (ours) 0.6M 0.6M 0.6M 0.6M 0.6M 5.6M 5.6M 14.2M
PointGPT-L [13] 360.5M 360.5M 360.5M 360.5M 360.5M 339.4M 339.3M 2481.2M
PointGST (ours) 2.4M 2.4M 2.4M 2.4M 2.4M 31.9M 31.8M 75.7M
TABLE XI: Comparison of different fine-tuning strategies on GPU consuming and inference speed (FPS). We report the results with a batch size of 8 on a single NVIDIA RTX 4090.
Methods Reference Params. (M) Infer. memory (G) FPS PB_T50_RS
PointGPT-L [13] NeurIPS 23 360.5 3.8 72.9 93.40
IDPT [16] ICCV 23 10.0 3.9 70.4 92.99(-0.41)
Point-PEFT [18] AAAI 24 3.1 4.0 32.4 92.85(-0.55)
DAPT [17] CVPR 24 4.2 3.9 66.7 93.02(-0.38)
PointGST (ours) - 2.4 3.9 71.3 94.83(+1.43)
TABLE XII: Inference time and GPU memory percentages for different components of our PointGST on ScanNetV2 [83] dataset (NVIDIA RTX 4090, batch size 1).
Metric Pre-processing Others Total
Graph construction Eigen-decomposition
Inference time (%) 0.42 1.60 97.98 100
GPU memory (%) 5.60 0.58 93.82 100
@@ 6.8 Visualization Results
6.9 Training and Inference Cost
Our method requires only fine-tuning the proposed lightweight Point Cloud Spectral Adapter (PCSA) while freezing the entire pre-trained backbone. For example, we just fine-tune about 0.67% of the parameters in the SOTA baseline, PointGPT-L [13], achieving promising results. As the number of downstream tasks or datasets requiring fine-tuning increases, we only need to store the task/dataset-specific PCSA and one original pre-trained model. As shown in Tab. X, when fine-tuning various downstream tasks, our PointGST significantly compresses model size compared to the baselines [8, 13].
We further analyze the training GPU memory, an important metric for evaluating the training cost. As shown in Fig. 7, the proposed PointGST is highly efficient, consuming considerably less GPU memory than the fully fine-tuned models. Even compared to other PEFT methods such as IDPT [16], DAPT [17], and Point-PEFT [18], our method performs better in terms of memory efficiency. Notably, methods like Point-PEFT and IDPT cause significant memory consumption spikes. For example, Point-PEFT experiences a substantial memory increase when scaling batch size, which may be problematic for GPUs with limited memory (e.g., 24 GB cards). During the testing phase, as shown in Tab. XI, while the inference speed (FPS) of PEFT methods, including ours, is slightly lower than that of the baseline (FFT), our method demonstrates the best FPS among all PEFT approaches. This is expected, as adapter-tuning-based methods typically freeze the pre-trained backbone and introduce lightweight learnable modules, which slightly impact inference speed due to additional operations. Besides, the inference GPU memory of our method is highly competitive with the FFT and other PEFT methods.
We also provide a detailed analysis of both inference time and GPU memory usage percentages for different components of PointGST on the large-scale ScanNetV2 dataset, with PointGPT-L [13] as the baseline, as shown in Tab. XII. We observe that the percentages of time and GPU memory consumed by graph construction and Laplacian eigen-decomposition constitute only a small fraction of the total inference process, with the vast majority allocated to other standard operations. These results indicate that the additional spectral-domain computations introduced by our method incur negligible overhead in both inference time and GPU memory usage.
TABLE XIII: The effect of each component of our PointGST. The trainable parameters (Params.) and the overall accuracy (%) are reported. The results before and after “/” denote using Point-MAE and PointGPT-L as baselines, respectively.
@@ 6.9 Training and Inference Cost
- - - 0.62/2.43 81.99/93.88
- - 0.62/2.43 83.03/94.27
- - 0.62/2.43 84.07/94.48
- 0.62/2.43 84.52/94.69
- - - 0.62/2.43 81.99/94.17
- - 0.62/2.43 83.38/94.45
- - 0.62/2.43 83.69/94.55
- 0.62/2.43 83.83/94.62
0.62/2.43 85.29/94.83
TABLE XIV: The comparison of different tuning types. The results before and after “/” denote using Point-MAE and PointGPT-L as baselines, respectively.
Tuning type Method Params. (M) PB_T50_RS
Spatial tuning Fully fine-tuning 22.1/360.5 85.18/93.40
IDPT [16] 1.7/10.0 84.94/92.99
DAPT [17] 1.1/4.2 85.08/93.02
PointGST w/o spectral 0.6/2.4 84.04/93.56
Spectral tuning PointGST (ours) 0.6/2.4 85.29/94.83
TABLE XV: Ablation study on the Point Cloud Spectral Adapter (PCSA), including dimension $r$, scale $s$, and the effect of different adaption choices. The trainable parameters (Params.) and the overall accuracy (%) are reported. The results before and after “/” denote using Point-MAE and PointGPT-L as baselines, respectively.
@@
6.10 Ablation Study
Following previous methods [17, 16], we conduct ablation studies on the challenging ScanObjectNN PB_T50_RS dataset. Unless otherwise specified, we choose Point-MAE [8] and PointGPT-L [13] as the baseline, and all experiments are performed without the voting strategy.
6.10.1 The effect of each component
We systematically evaluate the effect of each component within our PointGST by analyzing four key aspects: global spectral, local spectral, residual connections, and spectral inversion. The results are detailed in Tab. XIII.
Taking the Point-MAE baseline as an example, without any of our components, the linear probing setting achieves only 75.99% overall accuracy (OA), significantly lower than the fully fine-tuning counterpart by 9.19%. When incorporating either global or local spectral components individually, we observe a marked improvement of 6.00% in OA for both, while introducing only 0.35M additional trainable parameters. Further performance gains are seen when integrating residual connections and applying the inverse Graph Fourier Transform (iGFT) to project back into the spatial domain, aligning with the output of the transformer layer. Ultimately, the combination of all components leads to a significant increase in OA to 85.29%. This outcome underscores the complementary nature of global and local spectral representations, which can effectively de-correlate the inner confusion of point tokens and provide task-related intrinsic information to guide the fine-tuning phase.
To make a more direct comparison between spatial tuning and spectral tuning, we provide additional results where our method does not transform tokens into the spectral domain but instead processes the original spatial tokens directly. As shown in Tab. XIV, we observe that the proposed method achieves better performance compared with the spatial domain-tuning methods, demonstrating the importance of fine-tuning in the spectral domain. It also highlights the benefits of token de-correlation and intrinsic task-specific information provided by the spectral basis.
6.10.2 Ablation study on the hyper-parameters and design choices of PCSA
The effect of dimension $r$. This parameter is used to balance efficiency and performance. As shown in Tab. XV(a), as the dimension $r$ increases from 12 to 36, the performance of our method improves, with performance increasing from 83.66 to 85.29. However, further increasing $r$ to 96 slightly hurts the performance, with performance dropping to 84.98. Therefore, we set the dimension $r$ to 36 by default, which reports ideal performance and less trainable parameters.
The effect of scale $s$. This parameter controls the impact of spectral tokens on the task-agnostic tokens. As shown in Tab. XV(b), we empirically find that when the value is simply set as 111 without the need for extensive hyper-parameter tuning, the proposed method can achieve satisfactory performance.
The effect of different adaption choices. In our PCSA, we utilize a simple shared linear layer to adapt the global and local spectral tokens, shown in Fig. 2(b). Here, we explore other possible learnable choices, including shared Attention, MLP (two layers), DWConv, linear, and independent linear layers for global and local spectral. Note that for fair comparisons, we make these options meet similar trainable parameters (about 0.6M). The results presented in Tab. XV(c) validate that the used shared linear layer is simple yet effective compared with other learnable architectures.
TABLE XVI: The effect of inserted position. The trainable parameters (Params.) and the overall accuracy (%) are reported. The results before and after “/” denote using Point-MAE and PointGPT-L as baselines, respectively.
@@ 6.10.3 The effect of the position of PCSA
We then study the effect of the position of the PCSA. For the inserted layers, we consider the shallow layers and deep layers. Tab. XVI(a) clearly demonstrates that the performance of PointGST has a positive relative to the added layers. Taking the Point-MAE baseline as an example, injecting the PCSA into half of the layers will cause about a 2.12%/1.53% performance drop. Therefore, in our method, we inject the proposed point spectral adapter into all transformer layers. In addition, we also explore the effect of inserted modules. There are three potential choices: FFN, Attention, and each transformer block. As shown in Tab. XVI(b), inserting the PCSA into the FFN outperforms Attention and the entire transformer block by 1.01% and 1.04% overall accuracy for Point-MAE, respectively.
@@ 6.10.4 Analysis on the point cloud graph construction
The effect of different weighting schemes. To calculate the edge weight between two points in a graph, we employ a data-dependent scaling strategy. Here, we study various weighting schemes, as detailed in Tab. XVI(c). The most straightforward weighting scheme, Euclidean distance, yields sub-optimal performance with an overall accuracy of 85.15% and 94.73%. As discussed in Sec. 4.2.1, we contend that the closer two points are, the stronger the weight of their relationship should be. Therefore, we propose a data-dependent scaled distance for edge weight, which outperforms Euclidean distance by 0.14% and 0.10% on the Point-MAE and PointGPT-L baselines. Additionally, we evaluate cosine distance and Gaussian similarity [101] for each pair of key points, finding that our proposed data-dependent scaling strategy yields more stable results.
@@
The effect of different sorting methods and group number $k$. When constructing local sub-graphs, we sort and categorize the sampled key points into $k$ groups. This part evaluates the impact of various sorting methods, including random, k-nearest neighbors (KNN), and space-filling curves (e.g., Hilbert order and Z-order). Tab. XVI(d) shows that random sorting provides the least favorable results, while methods like KNN and space-filling curves that preserve spatial locality considerably enhance the performance. Among them, Trans Z-order sorting stands out as the most effective. KNN exhibits inferior performance, potentially because it does not yield stable results and sometimes aggregates identical points into different groups. In contrast, space-filling curves maintain a consistent scanning pattern, making it easier for models to learn from the data. Besides, as shown in Tab. XVI(e), we empirically find that the best performance is achieved when $k$ is set as 4.
Figure 8: Analysis on the point cloud graph.
(c) The effect of edge weights.
@@
Point cloud Euclidean Distance 85.15/94.73
@@
(d) The effect of sorting methods.
@@
Hilbert 84.00/94.45
Trans Hilbert 84.59/94.34
@@
(e) The effect of group number $k$.
@@
222 84.04/94.45
444 85.29/94.83
888 84.04/94.52
161616 84.84/94.55
Figure 9: Analysis on the spectral conversion.
@@
6.10.5 Analysis on Different Spectral Conversion
In this part, we investigate the effect of different spectral conversions, including the Fast Fourier Transform (FFT), Discrete Cosine Transform (DCT), and we used Graph Fourier Transform (GFT). As shown in Tab. 9, the results demonstrate the superiority of the GFT we applied. This can be attributed to the fact that the FFT and DCT are designed for regularly sampled or structured data and assume uniform spacing and regularity in the signals, making them unsuitable for processing non-Euclidean point clouds [23]. In contrast, to accommodate the irregular domain of point clouds, we apply the GFT, which is well-suited for such data structures and adapts to specific samples.
6.10.6 The Effect of the Input Point Number
Finally, to evaluate the effect of different input point densities, we conduct experiments on ModelNet40 [76] dataset using 128 (sparse), 1024 (default), and 8192 (dense) points as input, as summarized in Tab. XVII. PointGST consistently achieves the best or most competitive accuracy across all point regimes for both Point-MAE [8] and PointGPT-L [13] baselines. In particular, our method maintains high performance at the default 1024-point setting and exhibits strong robustness under extremely sparse (128) and dense (8192) inputs. For example, with the PointGPT-L baseline, our PointGST attains 91.7%, 94.8%, and 94.7% accuracy at 128, 1024, and 8192 points, respectively, outperforming all other parameter-efficient fine-tuning methods [16, 17, 18] in each case. This demonstrates that the proposed spectral adapter can effectively adapt to large variations in point cloud density, making our method suitable for diverse real-world scenarios where input may vary.
TABLE XVII: Classification accuracy on ModelNet40 [76] dataset with 128, 1024, and 8192 input points (no voting). Params. denotes the trainable parameters.
Pre-trained model Fine-tuning strategy Reference Params. (M) 128 1024 8192
Point-MAE [8] (ECCV 22) Fully fine-tune - 22.1 91.1 93.2 93.4
IDPT [16] ICCV 23 1.7 91.4 93.3 93.5
Point-PEFT [18] AAAI 24 0.7 91.1 93.3 93.6
DAPT [17] CVPR 24 1.1 91.2 93.5 93.2
PointGST (ours) - 0.6 91.5 93.5 93.8
PointGPT-L [13] (NeurIPS 23) Fully fine-tune - 360.5 91.5 94.1 92.7
IDPT [16] ICCV 23 10.0 89.3 93.4 92.0
Point-PEFT [18] AAAI 24 3.1 90.0 93.5 94.4
DAPT [17] CVPR 24 4.2 90.8 94.2 94.1
PointGST (ours) - 2.4 91.7 94.8 94.7
@@ 7 Conclusion
Limitation and Future Work. Currently, our spectral domain PEFT framework is not directly applicable to more challenging 3D generative tasks (e.g., text-to-3D or image-to-3D generation), which often require fundamentally different architectures and optimization strategies. Extending PointGST to such generative scenarios is a promising direction for future research.